# New Methods for Offline GstLAL Analyses

### v1

@@ II.1.2 IMBH template bank
In contrast to the AllSky bank, the IMBH bank is much smaller, and is only used to augment the GstLAL offline results in the IMBH parameter space. The IMBH bank covers the mass parameter space higher than what the AllSky template bank targets, as shown in Table 2. The lower limit of is set to such that the IMBH bank is an extention to the AllSky bank and the templates of the two template banks do not overlap in the - space, as shown in Fig. 1. The dimensionless spins extend from -0.7 to 0.99. The reason for the lower cutoff (-0.7) to be higher than that of the AllSky bank (-0.99) is to prevent the ringdown frequency of the templates from getting close to the lower frequency cutoff for matched filtering used by the IMBH search (10 Hz). Details of the parameters of the IMBH bank are shown in Tab. 2. Fig. 1 also shows a plot of the IMBH bank in the - space.
@@ III.7 New Extinction Model
The effect of applying the new extinction model on the noise triggers is shown in Fig. 6. We can see that in the noise regime (i.e. 50th percentile to the 99th percentile), the noise LR histogram with the new extinction model is very close to the candiate histogram. This shows that the new extinction model is successful in applying the effects of candidate clustering and relative contributions of template bins to the candidate set. By adopting the new extinction model, we get more accurate FARs, since the noise LR distribution used to assign FARs is more accurate than the old method. The new extinction model has also been adopted for GstLAL’s online operations in the second half of the fourth observing run (O4b).
@@ Acknowledgements.
This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/ ), a service of LIGO Laboratory, the LIGO Scientific Collaboration (LSC) and the Virgo Collaboration. We especially made heavy use of the LVK Algorithm Library. LIGO was constructed by the California Institute of Technology and the Massachusetts Institute of Technology with funding from the United States National Science Foundation (NSF) and operates under cooperative agreements PHYS- and PHY-. In addition, the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany supported the construction of Advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO) and construction and operation of the GEO600 detector. Additional support for aLIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. The authors are grateful for computational resources provided by the LIGO Lab culster at the LIGO Laboratory and supported by PHY- and PHY, the Pennsylvania State University’s Institute for Computational and Data Sciences gravitational-wave cluster, and supported by OAC-, PHY-, PHY-, OAC-, OAC-, and PHY-. The authors thank Thomas Dent (and the LVK CBC AllSky group) for sharing initial investigations on the Gaussian noise component of the noise model discussed in Sec. III.6. The authors also thank Reed Essick (and the entire LVK Rates and Populations group) for developing the injection sets used in the calculation of the weights for the AllSky and IMBH analyses. LT acknowledges NASA 80NSSC23M0104 and the Nevada Center for Astrophysics for support. CH Acknowledges generous support from the Eberly College of Science, the Department of Physics, the Institute for Gravitation and the Cosmos, the Institute for Computational and Data Sciences, and the Freed Early Career Professorship.References

### v2

@@ II.1.2 IMBH template bank
In contrast to the AllSky bank, the IMBH bank is much smaller, and is only used to augment the GstLAL offline results in the IMBH parameter space. The IMBH bank covers the mass parameter space higher than what the AllSky template bank targets, as shown in Table 2. The lower limit of is set to such that the IMBH bank is an extention to the AllSky bank and the templates of the two template banks do not overlap in the - space, as shown in Fig. 1. The dimensionless spins extend from -0.69 to 0.99. The reason for the lower cutoff (-0.69) to be higher than that of the AllSky bank (-0.99) is to prevent the ringdown frequency of the templates from getting close to the lower frequency cutoff for matched filtering used by the IMBH search (10 Hz). Details of the parameters of the IMBH bank are shown in Tab. 2. Fig. 1 also shows a plot of the IMBH bank in the - space.
@@ III.7 New Extinction Model
The effect of applying the new extinction model on the noise triggers is shown in Fig. 6. We can see that in the noise regime (i.e. 50th percentile to the 99th percentile), the noise LR histogram with the new extinction model is very close to the candiate histogram. This shows that the new extinction model is successful in applying the effects of candidate clustering and relative contributions of template bins to the candidate set. By adopting the new extinction model, we get more accurate FARs, since the noise LR distribution used to assign FARs is more accurate than the old method. The new extinction model has also been adopted for GstLAL’s online operations in the second part of the fourth observing run (O4b).
@@ Acknowledgements.
This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/ ), a service of LIGO Laboratory, the LIGO Scientific Collaboration (LSC) and the Virgo Collaboration. We especially made heavy use of the LVK Algorithm Library. LIGO was constructed by the California Institute of Technology and the Massachusetts Institute of Technology with funding from the United States National Science Foundation (NSF) and operates under cooperative agreements PHYS- and PHY-. In addition, the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany supported the construction of Advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO) and construction and operation of the GEO600 detector. Additional support for aLIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. The authors are grateful for computational resources provided by the LIGO Lab culster at the LIGO Laboratory and supported by PHY- and PHY, the Pennsylvania State University’s Institute for Computational and Data Sciences gravitational-wave cluster, and supported by OAC-, PHY-, PHY-, OAC-, OAC-, and PHY-. The authors thank Thomas Dent (and the LVK CBC AllSky group) for sharing initial investigations on the Gaussian noise component of the noise model discussed in Sec. III.6. The authors also thank Reed Essick (and the entire LVK Rates and Populations group) for developing the injection sets used in the calculation of the weights for the AllSky and IMBH analyses. LT acknowledges NASA 80NSSC23M0104 and the Nevada Center for Astrophysics for support. CH Acknowledges generous support from the Eberly College of Science, the Department of Physics, the Institute for Gravitation and the Cosmos, the Institute for Computational and Data Sciences, and the Freed Early Career Professorship. US and SS acknowledge support from NSF PHY-2409714.References